OpenStudy (anonymous):

Given the parent functions f(x) = 3x + 2 and g(x) = 5x − 10, what is g(x) − f(x)?

OpenStudy (anonymous):

The x after the 5 is small and looks like an exponent.

OpenStudy (anonymous):

the 2nd choice is correct. g(x)-f(x)= 5x-10-(3x+2) =5x-10-3x-2 =5x-10-2-3x =5x-12-3x Hope this helps!

OpenStudy (anonymous):

Ok, thanks:) Can you help me with one more, please?

OpenStudy (anonymous):

yes i will try

OpenStudy (anonymous):

Ok, thanks! If f(x) = log2 (x + 4), what is f−1(3)? *The 2 after log is small and lower than the log. The -1 small and to the top of f, like an exponent.

OpenStudy (anonymous):

3=log2x+4 solve for x "Inverse" may mean reciprocal, as so x^-1 = 1/x but it may mean going from ln x to exp(x), for example. Something is wrong, as none of your options correspond to my thought.

OpenStudy (anonymous):

Oh,sorry, I didn't give any options. These are them: 0 2 4 8

OpenStudy (anonymous):

Another, more roundabout, approach would be to actually find the inverse of y = base2log (x+4): a) replace f(x) with y b) interchange x and y (replace x with y and y with x) c) solve the resulting equation for y d) replace this y with -1 f (x) -1 e) The result is f (x) = 2^x -4.. f) evaluate this at x=3.

OpenStudy (anonymous):

I'm confused, so which answer would that be?

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

Wait, could it be 4? Substituting 3 for x? Idk

OpenStudy (anonymous):

i believe that it is 4 but it could be 2

OpenStudy (anonymous):

Well how did you solve it?

OpenStudy (anonymous):

i knew the answer to the first one but the 2nd one i used google for

OpenStudy (anonymous):

i would say 4 as the answer